OpenStudy (anonymous):

-4y + 17 less than or greater to 29 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution.

OpenStudy (ivettef365):

-4y + 17 <= 29 first subtract 17 on both sides, what do you get

OpenStudy (anonymous):

-4y<17

OpenStudy (ivettef365):

-4y + 17 <= 29 -17 -17 -4y <= 12

OpenStudy (anonymous):

thx. how describe?

OpenStudy (ivettef365):

now divide by -4, when you divide by a - number you need to switch the sign

OpenStudy (ivettef365):

what do you get

OpenStudy (anonymous):

-4y<-3?

OpenStudy (ivettef365):

the -4 next to the y cancels because -4/-4 = 1 so you have y => -3

OpenStudy (ivettef365):

so how does the graph goes is it an horizontal line or vertical line